OpenStudy (anonymous):

if (5 to the power of a)(5 to the power of b)=5 to the power of c/5 to the power of d, what is d in terms of a,b, and c?

hartnn (hartnn):

\(5^a.5^b=5^c/5^d\) any ideas how to start ?

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

a=2

OpenStudy (anonymous):

b=3

OpenStudy (anonymous):

c=5

OpenStudy (anonymous):

d=1

hartnn (hartnn):

that was yes/no question. and no,no

OpenStudy (anonymous):

why?

hartnn (hartnn):

\(\huge x^y.x^z=x^{y+z} \\ \huge \frac{x^y}{x^z}=x^{y-z}\) now try...

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

so

OpenStudy (anonymous):

look

OpenStudy (anonymous):

a=2

OpenStudy (anonymous):

b=3

OpenStudy (anonymous):

c=6!

OpenStudy (anonymous):

d=1

OpenStudy (anonymous):

choice B is the only answer

OpenStudy (anonymous):

6-2-3=1

hartnn (hartnn):

list your choices.

OpenStudy (anonymous):

c/ab

OpenStudy (anonymous):

c-a-b

OpenStudy (anonymous):

a+b-c

OpenStudy (anonymous):

c-ab

OpenStudy (anonymous):

and c/a+b

OpenStudy (anonymous):

choice b only works because 6-2-3=1

OpenStudy (anonymous):

the others do not arrive to d=1

hartnn (hartnn):

so, its choice B.....

OpenStudy (anonymous):

the correct answer!!!

OpenStudy (anonymous):

i am posting another problem now :)